# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

### v2

@@ III Simulation results
In the dynamic case, we study the performance of the approaches in a time-varying environment, where the approaches are implemented online to track the time-varying optimizer of (1) without waiting for them to converge. Figure 5 shows significant voltage limit violations without control and demonstrates that our proposed distributed nested feedback optimization approach successfully enforces the voltage limit even for the most remote, and therefore, the most sensitive bus. An average voltage violation (AVV), defined in Fig. 5, of $2.1\times 10^{-4}$ pu is achieved, which is only higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to single-point failures. The proposed nested feedback approach significantly outperforms other distributed ones, which is visible in Fig. 5 and quantified using the AVV.
IV Conclusion
@@
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed other distributed approaches.
@@
The proposed distributed approach does not extend to congestion management while the centralized does so. The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [1] to construct the sensitivity matrix $\mathbf{X}$. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach works for either active or reactive power control, but not joint control. This is because the observation in Section II-D does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.

### v3

@@ IV-C Dynamic case results
Table I shows the AVV values for bus 95 (the most sensitive bus) under various approaches. An AVV of $1.6\times 10^{-4}$ pu is achieved for the proposed approach, which is only higher than that of the centralized approach. The latter however requires a centralized communication architecture and is thus susceptible to single-point failures. The proposed nested feedback approach significantly outperforms other existing distributed or local approaches.
V Conclusion
@@
In this paper, we focused on the distributed voltage regulation problem in distribution systems and proposed a nested feedback optimization approach. We theoretically analyzed its convergence. Our simulation results showed that the approach, while only requiring short-range communication between physical neighbours, achieved satisfactory voltage regulation and outperformed existing distributed and local approaches.
@@
The proposed distributed approach does not extend to congestion management while the centralized does so. The current presentation is based on radial systems but the approach can be readily extended to meshed systems using the approach in [3] to construct the sensitivity matrix $\mathbf{X}$. Extending the approach to unbalanced systems is an interesting future direction, where the focus is to construct a positive definite $\mathbf{X}$ with a sparse inverse matrix. Finally, the approach works for either active or reactive power control, but not joint control. This is because the observation in Section III-B does not hold anymore since the voltage difference is no longer the gradient of either active or reactive power injection, but represents their joint non-separable impact. However, one might deploy active and reactive power feedback controllers separately at different sampling time. Whether the system under two feedback controllers is stable is left for future investigations.